# A Language Anchor-Guided Method for Robust Noisy Domain Generalization

### v1

@@ 4.2 Baseline Methods
Table 2: Cross-test accuracy (%) for domain shifts under a noise level of $\eta=$ 0.25. Best results in bold.
@@
Table 3: Cross-test accuracy (%) for domain shifts under a noise level of $\eta=$ 0.1. Best results in bold.
@@ 4.3 Experimental Results
Table 2 and 3 present the average test accuracy for different values of $\eta=0.1,0.25$ with each class in the dataset. From experimental results, $A^{3}W$ demonstrates consistent superior performance over all baseline methods. We observe that:
@@
3) Among the baselines, Mixup frequently attains the second-best performance, suggesting that interpolation-based augmentation helps mitigate moderate noise. However, these gains are relatively small and inconsistent across domains.
4) In our noise analysis (Figure 3), we observe the general tendency of an earlier decrease in accuracy during training as noise level increases due to noise overfitting. Among all evaluated methods, $A^{3}W$ exhibits the smallest decline in accuracy as noise rises, demonstrating that our method is less prone to overfitting under noisy conditions. Simple loss reweighting methods in noisy labels also have the advantage of having flexibility in adapting to network architecture.
In subsequent experiments, we use $A^{3}W$ as the default approach unless otherwise specified.
Figure 3: Noise analysis. (a) The effect of increasing noise levels on classification accuracy reveals that higher noise leads to a sharper decline in accuracy, reflecting an increased tendency to overfit. (b) Among the compared methods, $A^{3}W$ is more robust to noise injection, with its accuracy decreasing by only 0.2 when noise increases from 0.1 to 0.25, in contrast to other algorithms, which show declines between 0.427 and 0.527.
@@ 4.4 Ablation Study
To better understand the role of each component in $A^{3}W$, we conduct ablation studies under three key questions, as summarized in Table 4. The experiments span multiple datasets, and the results highlight how removing NLP anchor alignment or omitting the softmax weighting mechanism impacts overall performance.
(1) Why does removing NLP anchor alignment reduce performance?
In Table 4, discarding the NLP anchor alignment step (i.e., “w/out NLP anchor”) yields an average accuracy drop of about 0.86% across the listed datasets. This decrease arises because anchor alignment provides semantic guidance that helps the model differentiate meaningful features from spurious correlations. Without alignment, the feature extractor lacks an external semantic reference, making it more prone to overfitting on noisy labels and domain-specific artifacts.
(2) How does eliminating softmax weights affect training stability?
When we replace the adaptive softmax weighting with uniform weights (i.e., “w/out weighted loss”), the accuracy declines by an average of 1.16% across the datasets. This finding suggests that the continuous weighting scheme (Section 3.7) is crucial for emphasizing well-aligned (likely clean) samples. In contrast, uniform weighting fails to down-weight noisy or misaligned samples, reducing training stability and overall performance.
(3) Are both alignment and weighting necessary for robust generalization?
Finally, the baseline $A^{3}W$ model—which integrates both NLP anchor alignment and softmax weighting—achieves the highest accuracy on average (e.g., 80.76% on PACS). These results indicate that semantic anchors and adaptive weighting reinforce each other: alignment ensures meaningful feature extraction, while the weighting mechanism selectively highlights reliable data. Removing either component leads to noticeable performance degradation, underscoring their combined importance for robust domain generalization under noisy labels.
Table 4: Cross-test accuracy (%) for domain shifts. Best results in bold.
@@
Figure 5: t-SNE embeddings of three methods on the PACS dataset. From left to right are ERM, Mixup, and $A^{3}W$.
@@ 4.7 Impact of Semantic Richness in NLP Anchors
The NLP anchor has shown effective guidance in our ablation studies, which is further validated by experimental results on the SVIRO dataset, as shown in Table 2 and 3. In particular, when descriptive class names (e.g., “car back seat”, “infant car seat”, “child in convertible car seat”, etc.) were provided, the performance improvement of $A^{3}W$ was the highest among all datasets. In contrast, other datasets used single-word labels for each class. However, this dependency on rich semantic information can be problematic when the categories are very similar or less common. We eliminated the possibility that larger number of classes was the limitation, as our experiments on the OfficeHome dataset, which contains 65 classes, also showed improved performance compared to the baselines. Future work could investigate more tailored network architectures for instance-specific domain generalization, particularly for closely related classes.
@@ 4.8 Training Noise as an Implicit Regularizer
We also observed under some parameter settings a higher level of noise in training can sometimes lead to improved accuracy as shown in Fig. 3. This counterintuitive phenomenon occurs because additional noise acts as a form of regularization, discouraging the model from overfitting to spurious correlations and uninformative details in the training distribution. Based on a study from previous work  [64], the injected noise forces the network to learn more robust and invariant features by smoothing the loss landscape and encouraging the discovery of flatter minima. As a result, even though the training data contains more noise, the model focuses on the underlying, generalizable patterns that are shared with the target domain. This leads to better generalization performance on unseen data, as the network is less sensitive to the idiosyncrasies of the noisy training samples.

### v2

@@ 4.2 Baseline Methods
TABLE II: Cross-test accuracy (%) for domain shifts under a noise level of $\eta=$ 0.25. Best results in bold.
@@
TABLE III: Cross-test accuracy (%) for domain shifts under a noise level of $\eta=$ 0.1. Best results in bold.
@@ 4.3 Experimental Results
Table II and III present the average test accuracy for different values of $\eta=0.1$ and $\eta=0.25$, respectively. The experimental results show that $A^{3}W$ consistently outperforms all baseline methods. We observe the following:
@@
3) Among the baselines, Mixup frequently attains the second-best performance, suggesting that interpolation-based augmentation can help mitigate moderate noise. However, its improvements are relatively small and are inconsistent across different domains.
4) In our noise analysis (see Figure 3), we observe that the general training accuracy tends to drop earlier as noise levels increase, a sign of overfitting to noise. Among all evaluated methods, $A^{3}W$ exhibits the smallest decline in accuracy with rising noise, demonstrating its resilience to overfitting under noisy conditions. Additionally, the flexibility of simple loss reweighting in handling noisy labels allows it to adapt well to various network architectures. Thus, in subsequent experiments, we use $A^{3}W$ as the default approach unless otherwise specified.
Figure 3: Noise analysis. (a) The effect of increasing noise levels on classification accuracy reveals that higher noise leads to a sharper decline, reflecting an increased tendency to overfit. (b) $A^{3}W$ is most robust to noise injection, with its accuracy decreasing by only 0.2 when noise increases from 0.1 to 0.25, in contrast to other algorithms, which show declines between 0.427 and 0.527.
@@ 4.4 Ablation Study
To better understand the role of each component in $A^{3}W$, we conduct ablation studies under three key questions, as summarized in Table IV. The experiments span multiple datasets, and the results highlight how removing NLP anchor alignment or omitting the softmax weighting mechanism impacts overall performance.
1) Why does removing NLP anchor alignment reduce performance? In Table IV, discarding the NLP anchor alignment step (i.e., “w/out NLP anchor”) yields an average accuracy drop of about 0.86% across the listed datasets. This decrease arises because anchor alignment provides semantic guidance that helps the model differentiate meaningful features from spurious correlations. Without alignment, the feature extractor lacks an external semantic reference, making it more prone to overfitting on noisy labels and domain-specific artifacts.
2) How does eliminating softmax weights affect training stability? When we replace the adaptive softmax weighting with uniform weights (i.e., “w/out weighted loss”), the accuracy declines by an average of 1.16% across the datasets. This finding suggests that the continuous weighting scheme (Section 3.8) is crucial for emphasizing well-aligned (likely clean) samples. In contrast, uniform weighting fails to down-weight noisy or misaligned samples, reducing training stability and overall performance.
3) Are both alignment and weighting necessary for robust generalization? Finally, the baseline $A^{3}W$ model—which integrates both NLP anchor alignment and softmax weighting—achieves the highest accuracy on average (e.g., 80.76% on PACS). These results indicate that semantic anchors and adaptive weighting reinforce each other: alignment ensures meaningful feature extraction, while the weighting mechanism selectively highlights reliable data. Removing either component leads to noticeable performance degradation, underscoring their combined importance for robust domain generalization under noisy labels.
TABLE IV: Cross-test accuracy (%) for domain shifts. Best results in bold.
@@
Figure 5: t-SNE embeddings of three methods on the PACS dataset. From left to right, up to down are ERM, IRM, Mixup, and $A^{3}W$.
@@ 4.7 Impact of Semantic Richness in NLP Anchors
Our ablation studies reveal that NLP anchors provide effective guidance, a finding further supported by experimental results on the SVIRO dataset (see Table II and Table III). In particular, when descriptive class names such as “car back seat,” “infant car seat,” and “child in convertible car seat” were used, $A^{3}W$ achieved the highest performance improvement among all datasets. In contrast, other datasets used single-word labels showed less pronounced gains. However, this reliance on rich semantic information may pose challenges when the categories are very similar or less frequently encountered. We ruled out the possibility that the number of classes was a limiting factor—our experiments on the OfficeHome dataset, which includes 65 classes, still demonstrated performance gains over the baselines. Future work could explore more tailored network architectures for instance-specific domain generalization, especially in scenarios involving closely related classes.
@@ 4.8 Training Noise as an Implicit Regularizer
In some parameter settings, we observed that introducing more noise during training can actually lead to improved accuracy, as shown in Fig. 3. This counterintuitive result occurs because additional noise acts as a form of regularization. Essentially, it prevents the model from overfitting to spurious correlations and irrelevant details in the training data. Previous work [45] suggests that this injected noise encourages the network to learn more robust and invariant features by smoothing the loss landscape and promoting the discovery of flatter minima. As a result, even though the training data is noisier, the model is better able to capture the underlying patterns that generalize well to the target domain. This leads to improved performance on unseen data, as the network becomes less sensitive to the peculiarities of the noisy training samples.